A Bibliography of Publications of Reinhard Anton Muskens

Reinhard Anton Muskens

Tilburg University Department of Linguistics PO Box 90153 Tilburg

NL 5000 LE Holland

Tel: ?n/a? FAX: ?n/a?

E-mail: muskens@coli.uni-sb.de (Internet)

25 November 2011 Version 1.01

Abstract

This bibliography records publications of Reinhard Anton Muskens.

Title word cross-reference

Anaphora [Mus91a]. Attitudes [Mus93e].

Bathrooms [KM94].

Categorial [Mus94a]. Change [Mus91a, Mus95c]. Classical [Mus91b, Mus92, Mus97]. Combining [Mus96]. Compositional [Mus94b]. Computational [BMR94, BKMV97]. Coreference [Mus93a].

Discourse [Mus94b, KM95, Mus94a, Mus96]. Disjunction [KM95]. Dynamics [MVV97].

Existence [Mus93b].

First [BMR94]. Formulation [Mus89b].

Going [Mus89a]. Grained [Mus91b]. Grammar [Mus89a, Mus94a].

Hyperfine [Mus91b]. Hyperfine-Grained [Mus91b].

Independence [Mus95b]. Information [Mus93c]. International [BMR94, BKMV97].

Logic [Mus91a, Mus91b, Mus92, Mus95c, Mus97]. Logical [Mus92].

Meaning [Mus95a]. Meanings [Mus91b]. Modeltheoretic [Mus93b]. Montague [Mus89a, Mus96].

Negation [KM95].

Omniscience [Mus92]. Order [Mus95b]. Order-Independence [Mus95b].

Partial [Mus93c, Mus89a]. Partiality [Mus95a]. Perception [Mus93d]. Predicate [Mus93b]. Proceedings [BMR94, BKMV97]. Program [Mus97]. Propositional [Mus93e].

Relational [Mus89b]. Representation [KM95, Mus94b, Mus94a, Mus96].

Second [BKMV97]. Semantics [BMR94, BKMV97, Mus96, Mus97].

Tense [Mus95c]. Theory [KM95, Mus89b, Mus94b, Mus94a]. Types [Mus89b].

Umbrella [KM94]. Underspecification [Mus95b].

Verbs [Mus93d].

Workshop [BMR94, BKMV97].

References

Bunt:1997:PSI

[BKMV97] H. Bunt, L. Kievit, R. A.

Muskens, and M. Verlinden, editors. Proceedings of the Second International Workshop on Computational Semantics. Tilburg University, Tilburg, 1997.

Bunt:1994:PFI

[BMR94] H. Bunt, R. A. Muskens, and G. Rentier, editors. Proceedings of the First International Workshop on Computational Semantics. ITK, Tilburg University, Tilburg, 1994.

Krahmer:1994:UB

[KM94] E. Krahmer and R. A. Muskens. Umbrella’s and bathrooms. In M. Harvey and L. Santelmann, editors, Proceedings of SALT IV, pages 179–194. Cornell University, Cornell, 1994.

Krahmer:1995:NDD

[KM95] E. Krahmer and R. A. Muskens. Negation and Disjunction in Discourse Representation Theory. Journal of Semantics, 12:357–376, 1995.

Muskens:1989:GPM

[Mus89a] R. A. Muskens. Going Partial in Montague Grammar. In R. Bartsch, J. F. A. K. van Benthem, and P. van Emde Boas, editors, Semantics and Contextual Expression. Foris, Dordrecht, 1989.

Muskens:1989:RFT

[Mus89b] R. A. Muskens. A Relational Formulation of the Theory of Types. Linguistics and Philosophy, 12: 325–346, 1989.

Muskens:1991:ALC

[Mus91a] R. A. Muskens. Anaphora and the Logic of Change. In J. Van Eijck, editor, Logics in AI, Proceedings of JELIA ’90, pages 414– 430. Springer-Verlag, Berlin, Heidelberg, New York, Tokyo, 1991.

Muskens:1991:HGM

[Mus91b] R. A. Muskens. Hyperfine-Grained Meanings in Classical Logic. Logique et Analyse, 133/134:159–176, 1991. (Appeared in 1994).

Muskens:1992:LOC

[Mus92] R. A. Muskens. Logical Omniscience and Classical Logic. In D. Pearce and G. Wagner, editors, Logics in AI, Proceedings of Jelia ’92, pages 52–64. Springer-Verlag, Berlin, Heidelberg, New York, Tokyo, 1992.

Muskens:1993:C

[Mus93a] R. A. Muskens. Coreference. In R. E. Asher and J. M. Y. Simpson, editors, The Encyclopedia of Language and Linguistics, volume 2, page 769. Pergamon, New York, 1993.

Muskens:1993:EPM

[Mus93b] R. A. Muskens. Existence Predicate (Modeltheoretic). In R. E. Asher and J. M. Y. Simpson, editors, The Encyclopedia of Language and Linguistics, volume 3, page 1191. Pergamon, New York, 1993.

Muskens:1993:PI

[Mus93c] R. A. Muskens. Partial Information. In R. E. Asher and J. M. Y. Simpson, editors, The Encyclopedia of Language and Linguistics, volume 6, pages 2952–2953. Pergamon, New York, 1993.

Muskens:1993:PV

[Mus93d] R. A. Muskens. Perception Verbs.

In R. E. Asher and J. M. Y. Simpson, editors, The Encyclopedia of Language and Linguistics, volume 6, pages 2999–3000. Pergamon, New York, 1993.

Muskens:1993:PA

[Mus93e] R. A. Muskens. Propositional Attitudes. In R. E. Asher and J. M. Y. Simpson, editors, The Encyclopedia of Language and Linguistics, volume 6, pages 3380– 3383. Pergamon, New York, 1993.

Muskens:1994:CGD

[Mus94a] R. A. Muskens. Categorial Grammar and Discourse Representation Theory. In Proceedings of COLING 94: the 16th International Conference on Computational Linguistics, August 5–9, 1994, Kyoto, Japan, pages 508–

514. ICCL, Kyoto, Japan, 1994.

Muskens:1994:CDR

[Mus94b] R. A. Muskens. A Compositional Discourse Representation Theory. In P. Dekker and M. Stokhof, editors, Proceedings of the Ninth Amsterdam Colloquium, pages 467–

486. ILLC, University of Amster-

dam, Amsterdam, 1994.

Muskens:1995:MP

[Mus95a] R. A. Muskens. Meaning and Partiality. CSLI, Stanford, 1995.

Muskens:1995:OIU

[Mus95b] R. A. Muskens. Order-Independence and Underspecification. In J. Groenendijk, editor, Ellipsis, Underspecification, Events and More in Dynamic Semantics. DYANA Deliverable R.2.2.C,

                                        1995. ftp://illc-sun.illc.uva.nl/pub/dyana/R2.2.C/Muskens.ps.

Muskens:1995:TLC

[Mus95c] R. A. Muskens. Tense and the Logic of Change. In U. Egli, E. P. Pause, C. Schwarze, A. Von Stechow, and G. Wienold, editors, Lexical Knowledge in the Organization of Language, pages 147–183. John Benjamins, Amsterdam, 1995.

Muskens:1996:CMS

[Mus96] R. A. Muskens. Combining Montague Semantics and Discourse Representation. Linguistics and Philosophy, 19:143–186, 1996.

Muskens:1997:PSC

[Mus97] R. A. Muskens. Program Semantics and Classical Logic. CLAUS Report 86, Universität des Saarlandes, January 1997.

http://coli.uni-sb.de/claus/claus86.html.

Muskens:1997:D

[MVV97] R. A. Muskens, J. F. A. K. Van Benthem, and A. Visser. Dynamics. In J. F. A. K. Van Benthem and A. Ter Meulen, editors, Handbook of Logic and Language, pages 587–648. Elsevier Science Publishers, Amsterdam, The Netherlands, 1997.